### v2

# A Survey of Personalized Large Language Models: Progress and Future Directions

Abstract
Large Language Models (LLMs) excel in handling general knowledge tasks, yet they struggle with user-specific personalization, such as understanding individual emotions, writing styles, and preferences. Personalized Large Language Models (PLLMs) tackle these challenges by leveraging individual user data, such as user profiles, historical dialogues, content, and interactions, to deliver responses that are contextually relevant and tailored to each user’s specific needs. This is a highly valuable research topic, as PLLMs can significantly enhance user satisfaction and have broad applications in conversational agents, recommendation systems, emotion recognition, medical assistants, and more. This survey reviews recent advancements in PLLMs from three technical perspectives: prompting for personalized context (input level), finetuning for personalized adapters (model level), and alignment for personalized preferences (objective level). To provide deeper insights, we also discuss current limitations and outline several promising directions for future research.
Keywords Large Language Models Personalization Memory Text Generation Recommendation
Contents
1 Introduction
In recent years, substantial progress has been made in Large Language Models (LLMs) such as GPT, PaLM, LLaMA, DeepSeek, and their variants (Zhao et al., 2023). These models have demonstrated remarkable versatility, achieving state-of-the-art performance across various natural language processing (NLP) tasks, including question answering, logical reasoning, and machine translation (Chang et al., 2024; Hu et al., 2024; Zhang et al., 2024a, b; Zhu et al., 2024a; Wang et al., 2023a, 2024a; Li et al., 2024a; Zhao et al., 2025a; Zhu et al., 2025), with minimal task-specific adaptation.
The Necessity of Personalized LLMs (PLLMs)
While LLMs excel in general knowledge and multi-domain reasoning, their lack of personalization creates challenges in situations where user-specific understanding is crucial. For instance, conversational agents need to adapt to a user’s preferred tone and incorporate past interactions to deliver relevant, personalized responses. As LLMs evolve, integrating personalization capabilities has become a promising direction for advancing human-AI interaction across diverse domains such as recommendation, education, healthcare, and finance (Wang et al., 2024b; Hu et al., 2024; Zhang et al., 2024a, b; Zhu et al., 2024a; Wang et al., 2023a, 2024a). Despite its promise, personalizing LLMs presents several challenges. These include efficiently representing and integrating diverse user data, addressing privacy concerns, managing long-term user memories, inferring users’ implicit preferences, etc (Salemi et al., 2023). Moreover, achieving personalization often requires balancing accuracy and efficiency while addressing biases and maintaining fairness in the outputs.
Contributions
Despite growing interest, the field of PLLMs lacks a systematic review that consolidates recent advancements. This survey aims to bridge the gap by systematically organizing existing research on PLLMs and offering insights into their methodologies and future directions. The contributions of this survey are as follows:
A systematic formulation of personalization scenarios: We provide precise problem formulations for different personalized LLM scenarios based on query types and personalized data characteristics, establishing a unified framework that distinguishes between different personalized techniques (Section 2.2).
A structured taxonomy: We propose a comprehensive taxonomy, providing a technical perspective on the existing approaches to building PLLMs across three levels: input-level prompting, model-level adaptation, and objective-level alignment (Section 2.3).
- (3)
A benchmark and evaluation summarization: We provide summarization of metrics, benchmarks tailored to different query types (extraction, abstraction, and generalization) and personalization scenarios (Section 6).
Future directions: We highlight current limitations and outline promising avenues for future research, including multimodal personalization, edge computing, lifelong updating, trustworthiness, etc (Section 7).
2 Preliminary
2.1 Large Language Models
Large Language Models (LLMs) generally refer to models that utilize the Transformer architecture and are equipped with billions of parameters trained on trillions of text tokens. These models have demonstrated substantial improvements in a myriad of tasks related to natural language understanding and generation, increasingly proving beneficial in assisting human activities. In this work, we mainly focus on autoregressive LLMs, which are based on two main architectures: decoder-only models and encoder-decoder models. Encoder-decoder models such as Flan-T5 (Chung et al., 2022) and ChatGLM (Zeng et al., 2022) analyze input through the encoder for semantic representations, making them effective in language understanding in addition to generation. Decoder-only LLMs focus on left-to-right generation by predicting the next token in a sequence, with numerous instances (Brown et al., 2020; Chowdhery et al., 2022; Touvron et al., 2023; Guo et al., 2025) under this paradigm achieving breakthroughs in advanced capabilities.
However, these models are typically pre-trained on general-purpose data and lack an understanding of specific user information. As a result, they are unable to generate responses tailored to a user’s unique tastes and expectations, limiting their effectiveness in personalized applications where user-specific adaptation is critical.
2.2 Problem Statement
Personalized Large Language Models (PLLMs) generate responses that align with the user’s style and expectations, offering diverse answers to the same query for different users (Clarke et al., 2024).
Problem Formulation
Let be a finite set of users. For each :
, where and is the query space;
is the set of ’s personalized data, where denotes the space of all user personalized data.
Let be the output space. For each user and query , there exists a desired output that aligns with ’s preferences and expectations. We define a metric function that evaluates the quality of the LLM-generated output compared with the desired output (the summarization of metrics is shown in Section 6.1, Table 5). Here, we generally assume quantifies the degree of match between the desired and model outputs, with higher values indicating better personalization performance. All notations are summarized in Table 8.
Note that represents the tunable parameters that can exist either within the base LLM architecture itself (when fine-tuning is permitted) or within extra modules incorporated into the LLM framework, like the LoRA module (Yang et al., 2024). Based on personalized data , queries , and outputs (responses) , personalization can be further categorized into multiple scenarios (Figure 2).
Personalized Data
The personalized data may encapsulate information about the user’s preferences, history, context, and other user-specific attributes. These can include profile/relationship, historical dialogues, historical content, and predefined human preferences. A detailed description, including examples, of the classification of input data types for user is provided:
Profile/Relationship User profile, including attributes (e.g., name, gender, occupation), and relationships (e.g., friends, family members), such as .
Historical Dialogues Historical dialogues, such as question-answer pairs that user interacts with the LLM (e.g., ), where each is a query and is the answer.
Historical Content Includes documents, previous reviews, comments or feedback from user . For example, .
Historical Interactions Includes historical interactions, preferences, ratings from user . For example, .
Pre-defined Human Preference: Define a set containing of preference dimensions such as “Helpfulness”. Choose various combinations of these dimensions, form individual preferences, and incorporate them as the instruction. For example, a preference prompt could be “Be harmless and helpful”.
Query Type
Different queries have different focal points regarding the expected responses , which have a certain influence on the method design (high-level comparison with examples is shown in Figure 4 and Table 1). We primarily classify them into the following categories, with further benchmark-based demonstrations provided in Section 6:
Extraction (explicit) This type refers to factual lookup, where the answer can be directly found in the user’s personalized data . The model acts as a retriever:
Here denotes the set of explicit factual tuples contained in . Importantly, extraction also covers simple statistical or aggregative operations (e.g., counting, aggregation) applied over the facts in .
Abstraction (implicit) The model applies a summarization mapping that condenses into a query-independent summary . This abstract profile captures long-term tendencies or preferences, enabling responses beyond explicit tuples:
Generalization (implicit) The model constructs a query-specific state , dynamically combining the user’s data with the current query. The response is then generated from a conditional distribution that also leverages external knowledge :
Unlike , which is global and query-independent, is dynamic and tailored to each query. External knowledge is necessary when alone is insufficient.
| Type | Depends on query? | Depends on external ? | Example |
| Extraction | ✓ | ✗ | “What is my occupation?” Singer; “How many coffees have I bought?” 5 |
| Abstraction | ✗ | ✗ | “Based on the movies I watched, what is my preference?” Sweet & Salty |
| Generalization | ✓ | ✓ | “Rewrite this sentence”; “Recommend a new movie I haven’t seen before” |
Downstream Task
By incorporating personalized data, PLLMs enhance traditional LLMs, improving response generation, recommendation, and classification tasks.
Generation Generation tasks typically involve representing a sequence of strings, such as generating answers for users based on their personalized data and questions or generating content according to the user’s writing style to assist their writing, and so forth (Salemi et al., 2023; Kumar et al., 2024; Zhao et al., 2025d; Au et al., 2025).
Note that our survey differs significantly from role-play related LLM personalization (Tseng et al., 2024; Chen et al., 2024b; Zhang et al., 2024f). While role-play focuses on mimicking characters during conversations (Zhu et al., 2024b; Zhao et al., 2024), PLLMs in this survey focus on understanding users’ contexts and preferences to meet their specific needs. Compared to (Zhang et al., 2024f), which emphasizes broad categories, our work provides a quite different systematic analysis of techniques to enhance PLLM efficiency and performance, with a detailed technical classification.
2.3 Proposed Taxonomy
We propose a taxonomy (as illustrated in Figure 2 and Figure 3) from technical perspectives (i.e., regarding the personalization operation ), categorizing the methods for Personalized Large Language Models (PLLMs) into three major levels:
Input level: Personalized Prompting focuses on handling user-specific data outside the LLM and injecting it into the model (Section 3).
Model level: Personalized Adaptation emphasizes designing a framework to efficiently fine-tune or adapt model parameters for personalization (Section 4).
Objective Level: Personalized Alignment aims to refine model behavior to align with user preferences effectively (Section 5).
3 Personalized Prompting
Prompt engineering acts as a bridge for interaction between users and LLMs. In this survey, prompting involves guiding an LLM to generate desired outputs using various techniques, from traditional text prompts to advanced methods like soft embedding. Soft embedding can be extended not only through input but also via cross-attention or by adjusting output logits, enabling more flexible and context-sensitive responses. For each user , the personalization operation is expressed as
where, is a function that extracts relevant context from the user’s personal context ; represents the combination operator that fuses the query and the relevant personalized context , producing enriched information for the LLM. Based on different designs of the operation, we further categorize prompting-based methods into four types, as shown in Figure 5 and Table 2: (1) profile-augmented prompting; (2) retrieval-augmented prompting; (3) soft-fused prompting; and (4) contrastive prompting.
3.1 Profile-Augmented Prompting
Profile-augmented prompting (Figure 5 (a)) explicitly utilizes summarized user preferences and profiles in natural language to augment LLMs’ input at the token level ( is the summarizer model). The summarizer model is typically distinct from the base LLM (generator).
Non-tuned Summarizer A frozen LLM can be directly used as the summarizer to summarize user profiles due to its strong language understanding capabilities, i.e., . For instance, Cue-CoT (Wang et al., 2023b) employs chain-of-thought prompting for personalized profile augmentation, using LLMs to extract and summarize user status (e.g., emotion, personality, and psychology) from historical dialogues. PAG (Richardson et al., 2023) leverages instruction-tuned LLMs to pre-summarize user profiles based on historical content. The summaries are stored offline, enabling efficient personalized response generation while meeting runtime constraints. ONCE (Liu et al., 2024a) prompts closed-source LLMs to summarize topics and regions of interest from users’ browsing history, enhancing personalized recommendations. DPL (Qiu et al., 2025a) enhances personalization through inter-user comparison, identifying similar users in the same cluster. Leveraging central users from the same cluster where the target user is located creates more accurate preference summaries.
Tuned Summarizer Black-box LLMs are sensitive to input noise, like off-topic summaries, and struggle to extract relevant information. Thus, training the summarizer to adapt to user preferences and style is essential. Matryoshka (Li et al., 2024b) uses a white-box LLM to summarize user histories, similar to PAG, but fine-tunes the summarizer instead of the generator LLM. RewriterSlRl (Li et al., 2024c) rewrites the query instead of concatenating summaries, optimized with supervised and reinforcement learning.
| Profile-Augmented (Section 3.1) | Retrieval-Augmented (Section 3.2) | Soft-Fused (Section 3.3) | Contrastive (Section 3.4) | |
| Summerizer () | Retriever | Encoder | Extractor | |
| Token Combination | Token Combination | Input Prefix; Attention; Logits | Hidden States; Logits | |
| Pros. | Efficiency | Long-Term Memory | Semantic Nuances |
Interpretability
Controllable |
| Cons. | Infromation Loss | Computational Limits Irrelevant Data | Lack of Interpretability | Hyperparameter Sensitivity |
3.2 Retrieval-Augmented Prompting
Retrieval-augmented prompting (Gao et al., 2023; Fan et al., 2024; Qiu et al., 2024) excels at extracting the most relevant records from user data to enhance PLLMs (See Figure 5 (b)). Due to the complexity and volume of user data, many methods use an additional memory for more effective retrieval (Tan et al., 2025). Common retrievers include sparse (e.g., BM25 (Robertson et al., 1995)), and dense retrievers (e.g., Faiss (Johnson et al., 2019), Contriever (Izacard et al., 2021)). These methods effectively manage the increasing volume of user data within the LLM’s context limit, improving relevance and personalization by integrating key evidence from the user’s personalized data.
3.2.1 Personalized Memory Construction
This part designs mechanisms for retaining and updating memory to enable efficient retrieval of relevant information.
Non-Parametric Memory This category maintains a token-based database, storing and retrieving information in its original tokenized form without using parameterized vector representations. For example, MemPrompt (Madaan et al., 2022) and TeachMe (Dalvi et al., 2022) maintain a dictionary-based feedback memory (key-value pairs of mistakes and user feedback). MemPrompt focuses on prompt-based improvements, whereas TeachMe emphasizes continual learning via dynamic memory that adapts over time. MaLP (Zhang et al., 2024c) further integrates multiple memory types, leveraging working memory for immediate processing, short-term memory (STM) for quick access, and long-term memory (LTM) for key knowledge. FERMI (Kim and Yang, 2024) maintains misaligned responses and creates personalized prompts for users by using LLMs to progressively refine prompting strategies based on user profiles and past opinions.
Parametric Memory Recent studies parameterize and project personalized user data into a learnable space, with parametric memory filtering out redundant context to reduce noise. For instance, LD-Agent (Li et al., 2024d) maintains memory with separate short-term and long-term banks, encoding long-term events as parametric vector representations refined by a tunable module and retrieved via an embedding-based mechanism. MemoRAG (Qian et al., 2024), in contrast, adopts a different approach by utilizing a lightweight LLM as memory to learn user-personalized data. Instead of maintaining a vector database for retrieval, it generates a series of tokens as a draft to further guide the retriever, offering a more dynamic and flexible method for retrieval augmentation. Inspired by cognitive theories of memory, PRIME (Zhang et al., 2025f) introduces episodic and semantic memory mechanisms to enhance the personalization capabilities of LLMs, and employs a slow-thinking strategy to further improve personalized reasoning.
3.2.2 Personalized Memory Retrieval
The key challenge in the personalized retriever design lies in selecting not only relevant but also representative personalized data for downstream tasks. LaMP (Salemi et al., 2023) investigates how retrieved personalized information affects the responses of large language models (LLMs) through two mechanisms: in-prompt augmentation (IPA) and fusion-in-decoder (FiD). PEARL (Mysore et al., 2023) and ROPG (Salemi et al., 2024) similarly aim to enhance the retriever using personalized generation-calibrated metrics, improving both the personalization and text quality of retrieved documents. Meanwhile, HYDRA (Zhuang et al., 2024) trains a reranker to prioritize the most relevant information additionally from top-retrieved historical records for enhanced personalization. RPM (Kim et al., 2025a) focuses on reasoning ability by extracting user-specific factors from user historical personalized data and creating annotated reasoning paths. At inference, it retrieves these examples to guide reasoning-aligned outputs, strengthening personalized responses. HYDRA and RPM primarily target black-box models where access to the base LLM parameters is restricted.
3.3 Soft-Fused Prompting
Soft prompting differs from profile-augmented prompting in that it compresses personalized data into soft embeddings rather than summarizing it into discrete tokens. These embeddings are generated by a user feature encoder .
In this survey, we generalize the concept of soft prompting, showing that soft embeddings can be integrated (combination operator ) not only through the input but also via cross-attention or by adjusting output logits, allowing for more flexible and context-sensitive responses (See Figure 5 (c)).
Input Prefix Soft prompting, used as an input prefix, focuses on the embedding level by concatenating the query embedding with the soft embedding, and is commonly applied in recommendation tasks. PPlug (Liu et al., 2024b) constructs a user-specific embedding for each individual by modeling their historical contexts using a lightweight plug-in user embedder module. This embedding is then attached to the task input. UEM (Doddapaneni et al., 2024) is a user embedding module (transformer network) that generates a soft prompt conditioned on the user’s personalized data. PERSOMA (Hebert et al., 2024) enhances UEM by employing resampling, selectively choosing a subset of user interactions based on relevance and importance. REGEN (Sayana et al., 2024) combines item embeddings from user-item interactions via collaborative filtering and item descriptions using a soft prompt adapter to generate contextually personalized responses. PeaPOD (Ramos et al., 2024) personalizes soft prompts by distilling user preferences into a limited set of learnable, dynamically weighted prompts. Unlike previously mentioned methods, which focus on directly embedding user interactions or resampling relevant data, PeaPOD adapts to user interests by weighting a shared set of prompts. ComMer (Zeldes et al., 2025) tunes an encoder to compress historical documents into embeddings, merges them through mean pooling, then feeds the result into a frozen base LLM, distinguishing it from methods focused on historical interaction data. DEP (Qiu et al., 2025b) further models inter-user differences in the latent space instead of merely considering the user’s own information.
Cross-Attention Cross-attention enables the model to process and integrate multiple input sources by allowing it to attend to personalized data and the query. User-LLM (Ning et al., 2024) uses an autoregressive user encoder to convert historical interactions into embeddings through self-supervised learning, which are then integrated via cross-attention. The system employs joint training to optimize both the retriever and generator for better performance. RECAP (Liu et al., 2023) utilizes a hierarchical transformer retriever designed for dialogue domains to fetch personalized information. This information is integrated into response generation via a context-aware prefix encoder, improving the model’s ability to generate personalized, contextually relevant responses.
Output Logits GSMN (Wu et al., 2021) retrieves relevant information from personalized data, encodes it into soft embeddings, and uses them in attention with the query vector. Afterward, the resulting embeddings are concatenated with the LLM-generated embeddings, modifying the final logits to produce more personalized and contextually relevant responses.
3.4 Contrastive Prompting
The key insight of contrastive prompting is to utilize two forward paths of LLM to generate contrast pairs with and without personalized information (See Figure 5 (d)). By comparing model states with and without personalization, it identifies factors influencing personalization, enabling dynamic adjustment of the model’s personalization level. Two mainstream model states are utilized: hidden states (representations) and logits (predictive distributions).
CoS (He et al., 2024) is a special case that assumes there is a brief user profile for each query; it amplifies its influence in LLM response generation by combining output probabilities (logits) with and without the profile, i.e., , adjusting personalization degree through hyperparameter without fine-tuning. In contrast, StyleVector (Zhang et al., 2025b) obtains encoded personalized information in LLM via hidden representations. Specifically, given a user and the personalized data (user historical content), , it use GPT-3.5-Turbo to generate a general response for comparison with the expected personalized response , i.e., , where extracts hidden representations in base model ’s middle layer. extracts the user style-based vector, and by adding this vector with a scaling factor , this method can steer the LLM toward personalized output to a controllable degree. CoSteer (Lv et al., 2025) performs decoding-time personalization through local delta steering. A local SLM produces logits with and without personal context, and their difference is used to steer the cloud LLM’s logits iteratively, preserving privacy while achieving controllable personalization without fine-tuning.
3.5 Discussions
The three prompting methods have distinct pros and cons (Table 2):
This approach centers on maintaining, selecting, and integrating external personalized data into LLMs to improve user behavior comprehension. Overall, prompting-based methods are efficient and adaptable, enabling dynamic personalization with minimal computational overhead. Especially for RAG, which reduces LLM hallucination by grounding responses in retrieved evidence. Therefore, prompting methods are particularly suitable for explicit queries that require answering specific factual information, as mentioned in Section 2.2, excelling when personalization involves retrieving concrete information from user data.
However, they lack deeper personalization analysis, as they rely on predefined prompt structures to inject user-specific information and are limited in accessing global knowledge due to the narrow scope of prompts, which usually fail in tasks with implicit generation queries compared with adaptation and alignment methods (Tan et al., 2024a).
4 Personalized Adaptation
| One PEFT for All Users (Shared Parameters, Section 4.1) | One PEFT Per User (Individual Parameters, Section 4.2) | |
| Training Objective | ||
| Inference Stage | ||
| Pros. | Parameter Efficiency Scalability | Strong Personalization User Isolation |
| Cons. | Limited Personalization Persnonalized Data Dependency | Storage Overhead Training Complexity |
PLLMs require balancing fine-tuning’s deep adaptability with the efficiency of prompting. Therefore, specialized methods need to be specifically designed for PLLMs to address these challenges utilizing parameter-efficient fine-tuning methods (PEFT), such as LoRA (Hu et al., 2021; Yang et al., 2024), prefix-tuning (Li and Liang, 2021), MeZo (Malladi et al., 2023), etc. From an architectural perspective, these methods can be categorized into two types: (1) one PEFT for all users, where all users share a single fine-tuning module that differentiates personalized information, and (2) one PEFT per user, where each user has their own unique fine-tuning block. Details are illustrated in Figure 6 and Table 3.
4.1 One PEFT All Users
This method trains on all users’ data using a shared PEFT module, eliminating the need for separate modules per user. The shared module’s architecture can be further categorized.
Single PEFT PLoRA (Zhang et al., 2024d) and LM-P (Woźniak et al., 2024) utilize LoRA for PEFT of LLM, injecting personalized information via user embeddings and user IDs, respectively. PLoRA is further extended and supports online training and prediction for cold-start scenarios. UserIdentifier (Mireshghallah et al., 2021) uses a static, non-trainable user identifier to condition the model on user-specific information, avoiding the need for trainable user-specific parameters and reducing training costs. Review-LLM (Peng et al., 2024a) aggregates users’ historical behaviors and ratings into prompts to guide sentiment and leverages LoRA for efficient fine-tuning. However, these methods rely on a single architecture with fixed configurations (e.g., hidden size, insertion layers), making them unable to store and activate diverse information for personalization (Zhou et al., 2024). To solve this problem, MiLP (Zhang et al., 2024e) utilizes a Bayesian optimization strategy to automatically identify the optimal configuration for applying multiple LoRA modules, enabling efficient and flexible personalization.
Mixture of Experts (MoE) Several methods use the LoRA module, but with a static configuration for all users. This lack of parameter personalization limits adaptability to user dynamics and preference shifts, potentially resulting in suboptimal performance (Cai et al., 2024). RecLoRA (Zhu et al., 2024a) addresses this limitation by maintaining a set of parallel, independent LoRA weights and employing a soft routing method to aggregate meta-LoRA weights, enabling more personalized and adaptive results. Similarly, iLoRA (Kong et al., 2024) creates a diverse set of experts (LoRA) to capture specific aspects of user preferences and generates dynamic expert participation weights to adapt to user-specific behaviors.
Shared PEFT methods rely on a centralized approach, where user-specific data is encoded into a shared adapter by centralized LLMs. This limits the model’s ability to provide deeply personalized experiences tailored to individual users. Furthermore, using a centralized model often requires users to share personal data with service providers, raising concerns about the storage, usage, and protection of this data.
4.2 One PEFT Per User
Equipping a user-specific PEFT module makes LLM deployment more personalized while preserving data privacy. However, the challenge lies in ensuring efficient operation in resource-limited environments, as users may lack sufficient local resources to perform fine tuning.
No Collaboration There is no collaboration or coordination between adapters or during the learning process for each use in this category. UserAdapter (Zhong et al., 2021) personalizes models through prefix-tuning, fine-tuning a unique prefix vector for each user while keeping the underlying transformer model shared and frozen. PocketLLM (Peng et al., 2024b) utilizes a derivative-free optimization approach, based on MeZo (Malladi et al., 2023), to fine-tune LLMs on memory-constrained mobile devices.
Collaborative Efforts The “one-PEFT-per-user" paradigm without collaboration is computationally and storage-intensive, particularly for large user bases. Additionally, individually owned PEFTs hinder community value, as personal models cannot easily share knowledge or benefit from collaborative improvements. OPPU (Tan et al., 2024a) frames the personalized LLMs training as a two-stage process, where a global LoRA module is learned to capture shared knowledge across users initially, followed by user-specific LoRA modules for individual adaptation. PerFit (Liu et al., 2025) further enhances this by replacing LoRA modules with representation fine-tuning modules, motivated by their finding that the personalization of LLM correlates with a low-rank collaborative shift and diverse personalized shifts following the collaborative one. CoPe (Bu et al., 2025) utilizes reward-guided decoding tailored for personalization, aiming to maximize the implicit reward signal for each user. PER-PCS (Tan et al., 2024b) enables efficient and collaborative PLLMs by sharing a small fraction of PEFT parameters across users. It first divides PEFT parameters into reusable pieces with routing gates and stores them in a shared pool. For each target user, pieces are autoregressively selected from other users, ensuring scalability, efficiency, and personalized adaptation without additional training. PROPER (Zhang et al., 2025c) introduces hierarchical user grouping with MoE-LoRA integration, enabling fine-grained personalization through staged adaptation from population-level, group-level, to individual-level models.
Another efficient collaborative strategy is based on the federated learning (FL) framework. For example, Wagner et al. (2024) introduces a FL framework for on-device LLM fine-tuning, using strategies to aggregate LoRA model parameters and handle data heterogeneity efficiently, outperforming purely local fine-tuning. FDLoRA (Qi et al., 2024) introduces a personalized FL framework using dual LoRA modules to capture personalized and global knowledge. It shares only global LoRA parameters with a central server and combines them via adaptive fusion, enhancing performance while minimizing communication and computing costs.
There are other frameworks that can be explored, such as HYDRA (Zhuang et al., 2024), which also employs a base model to learn shared knowledge. However, in contrast to federated learning, it assigns distinct heads to each individual user to extract personalized information.
4.3 Discussions
Fine-tuning methods enable deep personalization by modifying a large set of model parameters, and parameter-efficient fine-tuning methods (e.g., prefix vectors or adapters) reduce computational cost and memory requirements while maintaining high personalization levels. These methods improve task adaptation by tailoring models to specific user needs, enhancing performance in tasks like sentiment analysis and recommendations. They also offer flexibility, allowing user-specific adjustments while leveraging pre-trained knowledge. However, they still face several challenges:
Addressing these issues calls for complementary strategies such as federated learning, synthetic data generation, and hybrid prompting–adaptation pipelines, in order to balance efficiency, personalization depth, and reliability.
5 Personalized Alignment
Alignment techniques (Bai et al., 2022; Rafailov et al., 2024) typically optimize LLMs to match the generic preferences of humans. However, in reality, individuals may exhibit significant variations in their preferences for LLM responses across different dimensions like language style, knowledge depth, and values. Personalized alignment seeks to further align with individual users’ unique preferences beyond generic preferences. A significant challenge in personalized alignment is creating high-quality user-specific preference datasets, which are more complex than general alignment datasets due to data sparsity. The second challenge arises from the need to refine the canonical RLHF framework (Ouyang et al., 2022) to handle the diversification of user preferences, which is essential for integrating personalized preferences without compromising efficiency and performance.
5.1 Data Construction
High-quality data construction is critical for learning PLLMs, primarily involving self-generated data through interactions with the LLM. Wu et al. (Wu et al., 2024a) constructs a dataset for aligning LLMs with individual preferences by initially creating a diverse pool of 3,310 user personas, which are expanded through iterative self-generation and filtering. This method is similar to PLUM (Magister et al., 2024) that both simulate dynamic interactions through multi-turn conversation trees, allowing LLMs to infer and adapt to user preferences. To enable LLMs to adapt to individual user preferences without re-training, Lee et al. (Lee et al., 2024) utilizes diverse system messages as meta-instructions to guide the models’ behavior. To support this, the MULTIFACETED COLLECTION dataset is created, comprising 197k system messages that represent a wide range of user values. To facilitate real-time, privacy-preserving personalization on edge devices while addressing data privacy, limited storage, and minimal user disruption, Qin et al. (Qin et al., 2024) introduces a self-supervised method that efficiently selects and synthesizes essential user data, improving model adaptation with minimal user interaction.
Research efforts are also increasingly concentrating on developing datasets that assess models’ comprehension of personalized preferences. Kirk et al. (Kirk et al., 2024) introduces PRISM Alignment Dataset that maps the sociodemographics and preferences of 1,500 participants from 75 countries to their feedback in live interactions with 21 LLMs, focusing on subjective and multicultural perspectives on controversial topics. PersonalLLM (Zollo et al., 2024) introduces a novel personalized testdb, which curates open-ended prompts and multiple high-quality responses to simulate diverse latent preferences among users. It generates simulated user bases with varied preferences from pre-trained reward models, addressing the challenge of data sparsity in personalization. ALOE (Wu et al., 2024a) further generates tree-structured multi-turn conversations instead of single-turn pairwise responses.
5.2 Personalized Alignment Optimization
| Training-Time Personalization | Decoding-Time Personalization | |
| Training Stage | (for each in ) | |
| Inference Stage | (Merging) (Ensemble) | |
| Pros. | Strong Personalization Efficient Inference | High Flexibility No Retraining |
| Cons. | High Training Cost Lack Flexibility | Storage Overhead High Inference Cost |
Personalized preference alignment is usually modeled as a multi-objective reinforcement learning (MORL) problem, where personalized preference is determined as the user-specific combination of multi-preference dimensions. Based on this, a typical alignment paradigm is training-time personalization. This involves using a personalized reward derived from multiple reward models to guide the training of policy LLMs, which is illustrated on the left of Figure 7. MORLHF (Wu et al., 2023) separately trains reward models for each dimension and retrains the policy language models using proximal policy optimization, guided by a linear combination of these multiple reward models. This approach allows for the reuse of the standard RLHF pipeline. MODPO (Zhou et al., 2023) introduces a novel RL-free algorithm extending Direct Preference Optimization (DPO) for managing multiple alignment objectives. It integrates linear scalarization into the reward modeling process, enabling the training of LMs using a margin-based cross-entropy loss as implicit collective reward functions.
The alternative paradigm is decoding-time personalization, which combines multiple trained policy LLMs during inference. Table 4 provides a detailed comparison of this approach against training-time personalization. Its main strategies, personalized weight merging and personalized model ensemble, are illustrated in Figure 7(b) and (c), respectively. Personalized Soups (Jang et al., 2023) and Reward Soups (Rame et al., 2024) address the challenge of RL from personalized human feedback by first training multiple policy models with distinct preferences independently and then merging their parameters post-hoc during inference. Both methods allow for dynamic weighting of the networks based on user preferences, enhancing model alignment and reducing reward misspecification. Also, the personalized fusion of policy LLMs can be achieved not only through parameter merging but also through model ensembling. MOD (Shi et al., 2024) outputs the next token from a linear combination of all base models, allowing for precise control over different objectives by combining their predictions without the need for retraining. The method demonstrates significant effectiveness when compared to the parameter-merging baseline. PAD (Chen et al., 2024a) leverages a personalized reward modeling strategy to generate token-level rewards that guide the decoding process, enabling the dynamic adaptation of the base model’s predictions to individual preferences.
However, these approaches are limited by their reliance on predefined personalization dimensions and a small set of user preferences, which fail to capture the diversity of real-world personalization needs. Recent work has therefore sought to extend reward learning to better accommodate user-specific preferences. For example, VPL (Poddar et al., 2024) employs a variational encoder to encode a small number of preference annotations from a given user into a latent variable that captures individual taste. A reward model is then conditioned on the latent variable and trained by maximizing the evidence lower bound, enabling accurate prediction of rewards aligned with the user’s unique preferences. PREF (Shenfeld et al., 2025) models each user’s personalized reward function as a linear combination of shared base reward functions, which are first learned from multi-user data via matrix factorization. Using an active learning strategy, it can efficiently infer a new user’s combination weights with only a handful of feedback examples (e.g., 10), enabling user-specific adaptation of LLMs’ outputs without retraining. PPT (Lau et al., 2024) leverages in-context learning for scalable personalization by generating two candidate responses for each user prompt, soliciting user rankings, and dynamically incorporating this feedback into the model’s context to adapt to individual preferences over time. Drift (Kim et al., 2025b) proposes a training-free method where user preference is modeled as a linear combination of interpretable attributes, and then computes a reward signal by comparing log-likelihoods under attribute-specific prompts to guide the generation of a frozen base model. In contrast to these offline methods based on static preferences, other approaches focus on real-time adaptation. AMULET (Zhang et al., 2025g) approaches this from another unique test-time perspective, by formulating the decoding process of each token as an independent online learning problem. It obtains the optimization direction by contrasting the model’s output with and without user-provided prompts and utilizes an efficient closed-form solution for real-time iterative optimization, thus adapting to personalized user needs instantly without retraining the model. RLPA (Zhao et al., 2025b) simulates real users via a dynamically updated user portrait model, enabling the language model agent to engage in multi-turn interactive reinforcement learning within this simulated environment, thus continuously adapting its policy to evolving user preferences. PersonaAgent (Zhang et al., 2025a) extends personalization from text style to decision behaviors such as tool usage, guided by a dynamic Persona that directs the agent’s decisions. Using test-time user-preference alignment, it compares the agent’s responses with the user’s real actions in real time, using textual loss as feedback to optimize the Persona, which ensures personalization run through every step of the tool decision.
5.3 Discussions
Current mainstream personalized alignment technologies mainly model personalization as multi-objective reinforcement learning problems, where personalized user preferences are taken into account during the training phase of policy LLMs via canonical RLHF, or the decoding phase of policy LLM via parameter merging or model ensembling. Typically, these methods are limited to a small number (e.g., three) of predefined preference dimensions, represented through textual user preference prompts. However, in real-world scenarios, there could be a large number of personalized users, and their preference vectors may not be known, with only their interaction history accessed. Consequently, developing more realistic alignment benchmarks to effectively assess these techniques is a critical area for future research.
6 Metric and Benchmark
As mentioned in Section 2.2, LLM personalization can be categorized into multiple types based on different inputs, query types, and tasks. Currently, numerous benchmarks are available to help us validate personalization approaches. This section primarily organizes these benchmarks according to the aforementioned aspects. We first list the commonly used metrics for various personalization tasks (Section 6.1), then review the different benchmarks and the metrics they employ (Section 6.2).
6.1 Metric
We employ a comprehensive set of mainstream evaluation metrics to assess model performance across different tasks. The detailed specifications of these metrics are presented in Table 5. We summarize Accuracy (Acc), F1 Score (F1), Matthews Correlation Coefficient (MCC) (Chicco and Jurman, 2020), Mean Absolute Error (MAE), and Root Mean Squared Error (RMSE) for the classification task; ROUGE-1 (R-1) (Lin, 2004), BLEU (Papineni et al., 2002), ROUGE-L (R-L) (Lin, 2004), METEOR (MTR) (Banerjee and Lavie, 2005), SBERT (Reimers and Gurevych, 2019), LLM-as-Evaluator (LLM-E) (Gu et al., 2024) metrics to evaluate the quality of generated text for the generation task; and Hit Ratio (HR), Precision, Recall, and NDCG for the recommendation task.
| Metric | Task Type | Formulation | Notes |
| Accuracy (Acc) | C (M) | Overall correctness | |
| F1 | C (M) | Precision and recall | |
| MCC (Chicco and Jurman, 2020) | C (M) | Class imbalance | |
| MAE (Chai and Draxler, 2014) | C (O) | Outliers | |
| RMSE (Hodson, 2022) | C (O) | Large deviation penalty | |
| Perplexity (Serban et al., 2016) | G | Model’s uncertainty | |
| Rouge-1 (R-1) (Lin, 2004) | G | N-gram overlap | |
| BLEU (Papineni et al., 2002) | G | Short length penalty | |
| Rouge-L (R-L) (Lin, 2004) | G | Longest common subsequence | |
| METEOR (MTR) (Banerjee and Lavie, 2005) | G | Linguistic features | |
| SBERT (Reimers and Gurevych, 2019; Zhang et al., 2020) | G | Vector-based signal | |
| LLM-as-Evaluator (LLM-E) (Gu et al., 2024) | G | Automated evaluation | |
| EGISES (Vansh et al., 2023) | G | Personalization insensitivity | |
| P-Accuracy (Vansh et al., 2023) | G | Personalization-aware accuracy | |
| PerSEval (Dasgupta et al., 2024) | G | Personalization evaluation | |
| Hit Ratio (HR) | R | At least one hit | |
| Precision (Gunawardana and Shani, 2009) | R | Proportion of relevant items | |
| Recall (Gunawardana and Shani, 2009) | R | Coverage of user’s relevant items | |
| NDCG (Wang et al., 2013) | R | Position-aware relevance metric |
6.1.1 Classification Task
For the classification task, , , , and represent true positives, true negatives, false positives, and false negatives, respectively; and denote the ground truth and predicted values.
Multi-class & Binary Classification (M)
In multi-class and binary classification tasks, where labels are categorical without inherent order, standard evaluation metrics such as Accuracy and F1 score are commonly employed to assess the model’s ability to correctly predict class membership. MCC provides a more robust assessment under class imbalance.
Ordinal Classification (O)
For ordinal multi-class classification, where labels possess a natural order or ranking, performance metrics like MAE and RMSE are preferred, as they account for the magnitude of prediction errors relative to the true order, providing a more nuanced evaluation of model quality. RMSE reflects the square-root of the mean of squared errors, providing a magnitude-sensitive measure that penalizes larger deviations more heavily. MAE complements this by computing the average absolute error, offering robustness to outliers.
6.1.2 Generation Task
For the generation task, and represent the generated and reference sequences, and indicates the length of the longest common subsequence between and . and denote the number of unique words in the candidate and reference texts, respectively. In the Perplexity formula, is the -th token in the reference text, represents all tokens that come before position , and is the probability that the model assigns to token given the preceding tokens. In the BLEU formula, the brevity penalty is based on and , which are the total word counts (including repeated words) in the candidate and reference sentences, respectively. In the METEOR formula, represents the number of matched words between the candidate text and reference text (including exact matches, stem matches, and synonym matches). represents the number of "chunks" in the matching sequence, where a chunk is a contiguous sequence of matched words. A higher number of chunks indicates less fluent or more fragmented matching between texts.
Human-based Evaluation (Human-E)
Human-E is a process where human judges assess LLM outputs by rating them on accuracy, helpfulness, coherence, appropriateness, etc. In this approach, human evaluators review and rate responses based on criteria such as accuracy, helpfulness, coherence, and appropriateness. Human-E is considered the gold standard because humans can detect nuances in language, recognize factual errors, understand contextual appropriateness, and make holistic judgments that automated metrics often miss.
Conventional Evaluation
BLEU, ROUGE-1, ROUGE-L, and METEOR to measure lexical overlap between the generated and ground-truth responses. BLEU evaluates n-gram precision focusing on system output, whereas ROUGE-1 measures unigram recall relative to reference texts. ROUGE-L extends this by capturing the longest common subsequence structure. METEOR incorporates linguistic features, including stemming and synonym matching, for enhanced evaluation. For semantic-level assessment beyond surface-level patterns, SBERT-based cosine similarity (Li et al., 2025a) provides a vector-based measure of meaning preservation between generated and reference responses.
LLM-based Evaluation (LLM-E)
Human-based evaluation provides the most reliable assessment of LLM response quality, but obtaining statistically significant results is costly and time-intensive. Therefore, the “LLM-as-a-judge” framework (Gu et al., 2024) is proposed that uses LLM to automatically evaluate the quality and relevance of generated text. By prompting LLMs to score or compare outputs, it offers scalable, context-aware, and semantically rich assessments with minimal human input. This approach surpasses traditional metrics in flexibility but faces challenges like model bias and consistency. It represents a promising method for automated, nuanced evaluation. ExPerT (Salemi et al., 2025b) introduces an explainable reference-based evaluation framework specifically designed for style-based personalized text generation. It leverages an LLM to extract atomic aspects and their evidence from the generated and reference texts, match the aspects, and evaluate their alignment based on content and writing style – two key attributes in personalized text generation.
Personalization Degree Evaluation (Per-E)
The aforementioned metrics are unable to assess the personalization capability of summarization models. To address this limitation, EGISES (Vansh et al., 2023) is proposed as the first automatic measure for evaluating personalization in text summarization. The core Deviation function measures summary-level personalization by computing the ratio averaged across all user pairs, where and represent the Jensen-Shannon divergence-weighted differences between users’ expected summaries and model-generated summaries, respectively. The Deviation function measures personalization quality: for good personalization, for poor personalization, where the model fails to match user expectation differences. Building upon EGISES, P-Accuracy (Vansh et al., 2023) provides a personalization-aware accuracy measure that penalizes traditional accuracy scores based on the model’s personalization capability, , where is the sigmoid function, and control the penalty intensity and personalization emphasis, respectively. However, EGISES suffers from the personalization-accuracy paradox, where models can achieve high responsiveness but low accuracy, leading to poor user experience. PerSEval (Dasgupta et al., 2024) addresses this limitation by introducing the EDP that penalizes EGISES scores when accuracy drops, computed as , where DEGRESS measures responsiveness, and EDP incorporates accuracy-based penalties to ensure that personalization evaluation considers both user preference alignment and content quality.
6.1.3 Recommendation Task
For the recommendation task, denotes the relevance score of the item at rank position , which may be binary (0 or 1) or graded (multi-level) depending on the evaluation context. In the NDCG formula, is the number of relevant items (limited to items), represents the relevance values sorted in descending order (the ideal ordering).
Item Recommendation
Conversational Recommendation
Conversational recommendation systems commonly use Recall and NDCG as evaluation metrics to measure coverage and ranking quality, employing the "LLMs-as-a-judger" framework (Zhao et al., 2025c; Huang et al., 2024; Sayana et al., 2024). Additionally, an LLM-based user simulator—creating unique personas via zero-shot ChatGPT prompting and defining preferences using dataset attributes—is also used to assess whether outputs align with user preferences (Huang et al., 2024). Note that in this scenario, we still categorize it as a generation task (G), as the core objective remains text generation in context, while the input data contains interaction information.
| Benchmark | Personalized Data | Query | Task | Technique | Metric |
| MemoryBank (Xue et al., 2025) | User Profile Dialogues | Extraction | G | PT | LLM-E |
| PerLTQA (Du et al., 2024) | User Profile Dialogues | Extraction | G | PT (RAG) | LLM-E |
LoCoMo ❀
(Maharana et al., 2024) |
User Profile Dialogues | Abstraction Extraction Generalization | G | PT (RAG) | F1, Recall |
| LongMemEval (Wu et al., 2025) | Dialogues | Extraction | G | PT (RAG) | LLM-E |
MMRC ❀
(Xue et al., 2025) |
Dialogues | Extraction | G | PT | LLM-E, Human-E |
| IMPLEXCONV (Li et al., 2025b) | Dialogues | Extraction Generalization | G | PT (RAG) | LLM-E |
| PrefEval (Zhao et al., 2025d) | Dialogues | Generalization | G , C | PT (RAG) SFT | Acc, LLM-E |
| HiCUPID (Mok et al., 2025) | Preference Dialogues | Extraction | G | PT (RAG), SFT AL (DPO) | BLEU, R-L, LLM-E |
| MemBench (Tan et al., 2025) | Dialogues | Abstraction Extraction Generalization | G | PT (RAG) | Acc |
| PER-CHAT (Wu et al., 2021) | User Profile Content | Extraction | G | PT (Soft) | Human-E, Perplexity, BLEU |
| LaMP (Salemi et al., 2023) | Content | Abstraction Generalization | G , C | PT, AD*, AL* | Acc, F1, MAE, RMSE, R-1, R-L |
| LongLaMP (Kumar et al., 2024) | Content | Generalization | G | PT (RAG) | R-1, R-L, MTR |
| PEFT-U (Clarke et al., 2024) | Content | Abstraction | C | FT | Acc |
| pGraphRAG (Au et al., 2025) | Content | Abstraction Generalization | G , C | PT (RAG) | MAE, RMSE, R-1, R-L, MTR |
| LaMP-QA (Salemi and Zamani, 2025) | Content | Generalization | G | PT (PAG + RAG) | LLM-E |
| DPL (Qiu et al., 2025a) | Content | Generalization | G | PT (PAG) | R-1, R-L, BLEU, MTR, LLM-E |
| PERSONABench (Li et al., 2025a) | Content | Abstraction Generalization | G , C | PT (ICL) | All (excl. LLM-E, Per-E) |
| PRISM (Kirk et al., 2024) | Preference | Generalization | G | AL* | – |
| PersonalLLM (Zollo et al., 2024) | Preference | Generalization | G | PT (ICL) | LLM-E |
| ALOE (Wu et al., 2024a) | User Profile Preference | Generalization | G | AL (DPO) SFT | LLM-E |
| REGEN (Sayana et al., 2024) | Interactions | Generalization | G | PT (Soft) | R-L, BLEU, SBERT |
| PersonalWAB (Cai et al., 2025) | User Profile Interactions | Generalization | R | PT + AL (RAG + SFT / DPO) | Acc |
| RecBench+ (Huang et al., 2025) | User Profile Interactions | Extraction Generalization | R | – | Precision, Recall |
6.2 Benchmark
Benchmarks for LLM personalization can be categorized based on the types of input data they utilize, the nature of the queries they address, and the specific tasks they are designed to evaluate. In terms of the personalized data format, benchmarks can be classified into four main types: Historical Content, Dialogues, Interactions, and Pre-defined Human Preference. Sometimes the user profiles are also included as auxiliary information. The mainstream benchmarks for each data type are summarized in Table 6 111PERSONABench in 6 is the abbreviation for PERSONACONVBENCH (Li et al., 2025a).222Among fact-based query benchmarks (such as PerLTQA (Du et al., 2024), LongMemEval (Wu et al., 2025), and IMPLEXCONV Li et al. (2025b)), some include memory-related testing tasks, like retrieval. Since these are not the focus of this paper, which investigates techniques for injecting and utilizing personalized user data to guide LLMs toward better personalized performance, these tasks are not listed in the summary..
Dialogue-based Data
For dialogue-based data, the benchmarks typically focus on the ability of LLMs to extract and utilize personalized information (e.g., user characteristics, schedules, preferences) from multi-turn conversations. The challenging aspect here is to accurately capture user intent and context from the long-term dialogue history, which may involve complex interactions, implicit preferences, and enormous unrelated content (Li et al., 2025b; Zhao et al., 2025d; Mok et al., 2025). Therefore, the mainstream techniques for this type of data are prompting-based methods, which can effectively leverage the most decisive information from the dialogue history without requiring extensive model retraining. The query types in these benchmarks often include explicit extraction of personalized factual details. Naturally, the metrics used are usually accuracy-based, such as F1 score and accuracy, with LLM-E as a preprocessing step to calculate the exact matching scores between the generated and ground-truth responses (Xue et al., 2025; Wu et al., 2025; Tan et al., 2025).
Historical Content-based Data
For historical content-based data, the benchmarks often evaluate how well LLMs can incorporate and reason over user-specific documents, such as past tweets, articles, or other textual materials. The key challenge is to enable the model to understand and integrate this external knowledge into its responses, which may require sophisticated retrieval and comprehension capabilities (Salemi et al., 2023; Kumar et al., 2024). Benefiting from the purer text content provided in these benchmarks, techniques like SFT are more commonly employed to intrinsically adapt the model to the user’s content, while the prompting-based RAG and PAG methods are auxiliarily used to further retrieve and conclude the most relevant terms for fine-tuning (Au et al., 2025; Salemi and Zamani, 2025). More query types of abstraction and generalization are included in these benchmarks, which require the model to not only extract factual information but also adapt its style and tone to match user preferences (Salemi et al., 2023). Therefore, the metrics used are more diverse, including both accuracy-based metrics and generation quality metrics like ROUGE, BLEU, and METEOR, with LLM-E and Human-E as supplementary evaluations (Li et al., 2025a).
Preference-based Data
For preference-based data, the benchmarks typically assess how well LLMs can align their outputs with user-defined preferences, which may be explicitly stated or implicitly inferred from user behavior. The main challenge is to accurately model and incorporate these preferences into the generation process, which may involve complex reasoning and decision-making (Kirk et al., 2024; Wu et al., 2024a). Techniques like RLHF and DPO are often employed to fine-tune the model based on user feedback, while prompting-based methods like ICL are also used to guide the model’s behavior during inference. Targeting the subjective nature of preferences, the query types in these benchmarks often involve open-ended generation tasks, where the model must produce responses that are not only factually accurate (Mok et al., 2025; Zollo et al., 2024) but also align with user tastes and styles (Kirk et al., 2024). Consequently, the metrics used are primarily generation quality metrics, with LLM-E and Human-E as critical components to evaluate the alignment with user preferences.
Interaction-based Data
For interaction-based data, the benchmarks often focus on how well LLMs can leverage user interaction histories, such as clicks, likes, and other engagement signals, to personalize recommendations and responses. The key challenge is to effectively model user behavior and preferences from these interactions, which may be sparse and noisy (Sayana et al., 2024; Huang et al., 2025). Techniques like fine-tuning are commonly used to adapt the model to user interaction patterns, while prompting-based methods are also employed to retrieve relevant information based on past interactions (Cai et al., 2025). The query types in these benchmarks often involve generalized recommendation tasks, where the model must generate personalized suggestions based on user behavior. Therefore, the metrics used are typically recommendation quality metrics, such as Hit Ratio and NDCG. Sometimes, the generation quality metrics like BLEU and ROUGE are also used when the recommendation is presented in a conversational format.
6.3 Discussions
Although existing benchmarks cover a wide range of data types, query types, and tasks, there are still some limitations and gaps that need to be addressed in future research. To generalize to more realistic scenarios, future benchmarks should consider the following aspects:
7 Vision and Future Directions
7.1 Vision and Challenges
The vision of PLLMs (Figure 8) is to endow models with the abilities to remember, adapt, and evolve. Remembering means grounding responses in user-specific knowledge from both short- and long-term memory. Adapting highlights the capacity to abstract and generalize from personalized data, enabling reasoning and dynamic adjustment to user preferences. Evolving emphasizes continual learning, where models update personalization without forgetting, where models update personalization without forgetting, staying aligned with users’ shifting goals and their growth over time.
(good/strong, capability already well-demonstrated);
(poor/weak, capability not well-supported);
(open/under-explored, requiring further research). We further analyze the vision of PLLMs from the perspective of three key techniques: prompting, adaptation, and alignment. As shown in Figure 8, each technique contributes differently across the stages of extraction, abstraction/generalization, and evolution, and each faces distinct trade-offs in terms of efficacy, efficiency, and privacy.
Efficacy At the extraction task, prompting is effective and efficient, enabling PLLMs to recall and leverage factual or personalized information with minimal overhead (Tan et al., 2025). However, as tasks move toward abstraction and generalization, the limitations of prompting become evident, and adaptation/alignment methods play a more central role (Tan et al., 2024b; Bu et al., 2025). These techniques allow models to internalize user-specific preferences, improving personalization depth, but often come at higher computational and data requirements. Finally, at the evolution stage, where continual updating and lifelong learning are essential, performance remains largely under-explored, as models struggle to grow and adapt without catastrophic forgetting.
Efficiency
Efficiency is a constant constraint across all goals. Extraction via prompting is highly efficient, but as adaptation and alignment become necessary for abstraction, efficiency drops sharply due to the cost of fine-tuning or maintaining additional modules (Clarke et al., 2024). At the evolution stage, efficiency challenges are amplified, since continual updating requires frequent synchronization between local devices and cloud servers, incurring latency, storage, and energy costs. This highlights a tension where methods that improve personalization depth often conflict with the need for scalable, lightweight deployment.
Privacy
A key challenge is maintaining user privacy while still enabling personalization. Per-user PEFT strategies provide strong privacy guarantees but at the cost of reduced performance (Peng et al., 2024b). In contrast, collaborative or federated strategies can enhance performance by leveraging cross-user signals, but inevitably raise the risk of privacy leakage during communication or model aggregation (Zhuang et al., 2024). Parameter-sharing approaches offer a compromise by avoiding direct exposure of raw user data, yet they shift the bottleneck to communication and synchronization efficiency in edge–cloud coordination (Qi et al., 2024).
Overall Trade-offs Taken together, the figure illustrates a triangular trade-off: methods that strongly guarantee privacy tend to limit performance; those that maximize performance often compromise privacy; and strategies that balance the two introduce efficiency bottlenecks. While prompting-based extraction already achieves good efficacy and efficiency, abstraction and especially lifelong evolution remain under-explored. Addressing this “trilemma” requires new techniques that can simultaneously enhance performance, protect privacy, and scale efficiently—a direction that we identify as a key frontier for future research.
7.2 Future Directions
Despite advances in Personalized Large Language Models (PLLMs), significant challenges persist, particularly in technical improvements. Current methods effectively handle basic user preferences but struggle with complex, multi-source data, especially in multimodal contexts like images and audio. Efficiently updating models on resource-constrained edge devices is also crucial. Fine-tuning enhances personalization but can be resource-intensive and difficult to scale. Developing small, personalized models through techniques like quantization could address these issues.
1.
Complex User Data While current approaches effectively handle basic user preferences, processing complex, multi-source user data remains a significant challenge. For example, methods that use user relationships in graph-like structures are still limited to retrieval augmentation (Du et al., 2024). How to effectively leverage this complex user information to fine-tune LLM parameters remains a significant challenge. Most methods focus on text data, while personalized foundation models for multimodal data (e.g., images, videos, audio) remain underexplored, despite their significance for real-world deployment and applications (Wu et al., 2024c; Pi et al., 2024; Shen et al., 2024; Xu et al., 2025).
2.
Edge Computing A key challenge in edge computing is efficiently updating models on resource-constrained devices (e.g., phones), where storage and computational resources are limited. For example, fine-tuning offers deeper personalization but is resource-intensive and hard to scale, especially in real-time applications. Balancing resources with personalization needs is important. A potential solution is to build personalized small models (Lu et al., 2024) for edge devices, using techniques like quantization and distillation.
3.
Edge-Cloud Collaboration The deployment of PLLMs in real-world scenarios encounters significant challenges in edge-cloud computing environments. Current collaborative approaches often lack efficient synchronization between cloud and edge devices, highlighting the need to balance local computation and cloud processing (Tian et al., 2024).
4.
Efficient Adaptation to Model Updates Updating fine-tuned PEFT parameters for each user when base LLM parameters change poses a challenge due to high user data volume and limited resources. Retraining costs can be prohibitive. Future research should focus on efficient methods for updating user-specific parameters without complete retraining, such as incremental learning and transfer learning. Moreover, the emerging paradigm of small language models (SLMs) (Liu et al., 2024d) offers a promising direction, where lightweight models can be co-trained or specialized to handle personalization updates more efficiently while reducing memory and energy costs.
5.
Lifelong Updating Given the large variety of user behaviors, a key challenge is preventing catastrophic forgetting while ensuring the efficient update of long-term and short-term of memory. Future research could explore continual learning (Wu et al., 2024d) and knowledge editing (Wang et al., 2024c; Zhang et al., 2024g) to facilitate dynamic updates of user-specific information.
Privacy Privacy remains a critical concern, particularly regarding user privacy when generating personalized responses. As LLMs are not typically deployed locally, risks of privacy leakage arise. Future research should focus on privacy-preserving methods, such as federated learning and differential privacy (Yao et al., 2024; Liu et al., 2024e), to protect user data effectively while leveraging the model’s capabilities.
Trustworthiness Beyond performance and efficiency, ensuring trustworthiness is critical for the adoption of personalized LLMs. Future research should address issues of interpretability, enabling users to understand why a personalized response is generated and to what extent it relies on their data (He et al., 2024), like neuron selection (Zhao et al., 2025e). Another direction is improving fairness and bias mitigation, since personalization risks amplifying stereotypes or reinforcing undesirable patterns if user-specific data is skewed. Ultimately, developing standardized evaluation protocols and benchmarks for trustworthiness in the context of personalization will be key to building systems that users not only find useful but can also confidently rely on.
Applications Exploring domain-specific challenges and opportunities for personalization is essential to advance the practical adoption of PLLMs. For example, in healthcare, it promises tailored treatments but must strictly safeguard privacy (Zhang et al., 2024c). In education, adaptive tutoring is enabled, yet fairness and long-term modeling remain concerns (Zhao et al., 2025f). In creative industries, they support style-aware generation while risking overfitting and reduced diversity. In enterprise applications (Li et al., 2025c), personalization enhances productivity and customer support, but requires scalable, consistent, and governable solutions.
| Method | Personalized Data | Query | LLM (Generator) | Retriever | Prompting | Fine-tuning | Task |
| §3 Personalized Prompting | |||||||
| Cue-CoT (Wang et al., 2023b) | Dialogues | Genl. | ChatGLM-6B, BELLE-LLaMA-7B-2M, ChatGPT, Alpaca-7B, Vicuna-7B-v1.1 | Token | G | ||
| PAG (Richardson et al., 2023) ✫ | Content | Abs., Genl. | GPT-3.5, Flan-T5, Vicuna-13B | Token | G, C | ||
| Matryoshka (Li et al., 2024b) ✫ | Content | Abs., Genl. | GPT-4o-mini, GPT-3.5-Turbo | Token | G, C | ||
| RewriterSlRl (Li et al., 2024c) | Content | Genl. | PaLM-2 | Token | G | ||
| R2P (Luo et al., 2025) ✫ | Content | Abs., Genl. | LLaMA-3.1-8B-Instruct, DeepSeek-R1-Distill-LLaMA-8B | Token | G, C | ||
| DPL (Qiu et al., 2025a) | Content | Genl. | GTE-Qwen2-1.5B-Instruct | Token | G | ||
| CoS (He et al., 2024) | User Profile | Abs., Genl. | GPT-3.5, LLaMA-2-7B-Chat, T0pp, Mistral-7B-Instruct | Token | G, C | ||
| StyleVector (Zhang et al., 2025b) ✫✫ | Content | Genl. | LLaMA-2-7B | Token | G | ||
| MemPrompt (Madaan et al., 2022) | Content | Genl. | GPT-3 | Token | G | ||
| TeachMe (Dalvi et al., 2022) | Content | Ext., Genl. | GPT-3 | Token | G | ||
| MaLP (Zhang et al., 2024c) | Dialogues | Abs., Genl. | GPT-3.5, LLaMA-7B, LLaMA-13B | Token | (LoRA) | G, C | |
| MemoRAG (Qian et al., 2024) | Dialogues | Ext., Genl. | Qwen2-7B-Instruct, Mistral-7B-Instruct | Token | G | ||
| IPA (Salemi et al., 2023) ✫ | Content | Abs., Genl. | Flan-T5-base | Token | G, C | ||
| FiD (Salemi et al., 2023) ✫ | Content | Abs., Genl. | Flan-T5-base | Token | G, C | ||
| MSP (Zhong et al., 2022) | Dialogues | Genl. | DialoGPT | Token | G | ||
| AuthorPred (Li et al., 2023) | Content | Abs., Genl. | T5-11B | Token | G, C | ||
| PEARL (Mysore et al., 2023) | Content | Genl. | Davinci-003, GPT-3.5-Turbo | Token | G | ||
| ROPG (Salemi et al., 2024) ✫ | Content | Abs., Genl. | Flan-T5-XXL-11B | Token | G, C | ||
| HYDRA (Zhuang et al., 2024) ✫ | Content | Abs., Genl. | GPT-3.5-Turbo | Token | (Adaptor) | G, C | |
| PRM (Zhuang et al., 2024) ✫ | Content | Abs., Genl. | GPT-4o-mini | Token | G, C | ||
| PersonaAgent (Zhang et al., 2025a) ✫ | Content | Abs., Genl. | Claude-3.5-Sonnet | Token | C | ||
| UEM (Doddapaneni et al., 2024) | Interactions | Abs. | Flan-T5-base, Flan-T5-large | Embedding | C | ||
| PERSOMA (Hebert et al., 2024) | Interactions | Genl. | PaLM-2 | Embedding | (LoRA) | G | |
| REGEN (Sayana et al., 2024) | Interactions | Genl. | PaLM-2 | Embedding | G | ||
| PeaPOD (Ramos et al., 2024) | Interactions | Genl. | T5-small | Embedding | G, R | ||
| ComMer (Zeldes et al., 2025) ✫✫ | Content | Genl. | Gemma-2B | Embedding | G | ||
| PPlug (Liu et al., 2024b) ✫ | Content | Abs., Genl. | Flan-T5-XXL-11B | Embedding | G, C | ||
| User-LLM (Ning et al., 2024) | Interactions | Genl. | PaLM-2-XXS | Embedding | G, R | ||
| RECAP (Liu et al., 2023) | Dialogues | Genl. | DialoGPT | Embedding | G | ||
| GSMN (Wu et al., 2021) | User profile Comment | Genl. | DialoGPT | Embedding | G | ||
| §4 Personalized Adaptation | |||||||
| PLoRA (Zhang et al., 2024d) | User profile (ID) | Abs. | BERT, RoBERTa, Flan-T5 | LoRA | C | ||
| UserIdentifier (Mireshghallah et al., 2021) | User profile | Abs. | RoBERTa-base | - | C | ||
| LM-P (Woźniak et al., 2024) | User profile (ID) | Abs., Genl. | Mistral-7B, Flan-T5, Phi-2, StableLM, GPT-3.5, GPT-4 | LoRA | G, C | ||
| Review-LLM (Peng et al., 2024a) | Interactions | Genl. | GPT-3.5-Turbo, GPT-4o, LLaMA-3-8B | LoRA | G | ||
| MiLP (Zhang et al., 2024e) | Content Dialogues | Genl. | DialoGPT, RoBERTa, LLaMA-2-7B, LLaMA-2-13B | LoRA | G | ||
| RecLoRA (Zhu et al., 2024a) | Interactions | Genl. | Vicuna-7B | LoRA | R | ||
| iLoRA (Kong et al., 2024) | Interactions | Genl. | LLaMA-2-7B | LoRA | R | ||
| UserAdapter (Zhong et al., 2021) | Interactions | Abs., Genl. | RoBERTa-base | Prefix-tuning | R | ||
| PocketLLM (Peng et al., 2024b) | Content | - | RoBERTa-large, OPT-1.3B | MeZo | C | ||
| OPPU (Tan et al., 2024a) ✫ | Content | Abs., Genl. | LLaMA-2-7B | (o) | (o) | LoRA | G, C |
| CoPe (Bu et al., 2025) ✫✫ | Content | Genl. | Mistral-7B-Instruct-v0.3 | LoRA | G | ||
| PER-PCS (Tan et al., 2024b) ✫ | Content | Abs., Genl. | LLaMA-2-7B | (o) | (o) | LoRA | G, C |
| PROPER (Zhang et al., 2025c) ✫ | Content | Abs., Genl. | LLaMA-2-7B | LoRA | G, C | ||
| (Wagner et al., 2024) | Content | Abs. | GPT-2 | LoRA | C | ||
| FDLoRA (Qi et al., 2024) | Content | - | LLaMA-2-7B | LoRA | C | ||
| §5 Personalized Alignment | |||||||
| Wu et al. (2024a) | Dialogues | Genl. | Qwen2-7B-Instruct, LLaMA-3-8B-Instruct, Mistral-7B-Instruct-v0.3 | G | |||
| PLUM (Magister et al., 2024) | Dialogues | Genl. | LLaMA-3-8B-Instruct | LoRA | G | ||
| Lee et al. (2024) | User Profile | Genl. | Mistral-7B-v0.2 | G | |||
| MORLHF (Wu et al., 2023) | Preference | Genl. | GPT-2, T5-large | G | |||
| MODPO (Zhou et al., 2023) | Preference | Genl. | LLaMA-7B | LoRA | G | ||
| Personalized Soups (Jang et al., 2023) | Preference | Genl. | Tulu-7B | LoRA | G | ||
| Reward Soups (Rame et al., 2024) | Preference | Genl. | LLaMA-7B | LoRA | G | ||
| MOD (Shi et al., 2024) | Preference | Genl. | LLaMA-2-7B | G | |||
| PAD (Chen et al., 2024a) | Preference | Genl. | LLaMA-3-8B-Instruct, Mistral-7B-Instruct | LoRA | G | ||
| PPT (Lau et al., 2024) | Preference | Genl. | Self-defined | G | |||
| VPL (Poddar et al., 2024) | Preference | Genl. | GPT-2, LLaMA-2-7B | LoRA | G | ||
8 Conclusions
This survey offers a comprehensive overview of PLLMs, focusing on personalized responses to individual user data. It presents a taxonomy categorizing approaches into three key perspectives: Personalized Prompting (Input Level), Personalized Adaptation (Model Level), and Personalized Alignment (Objective Level), with further subdivisions. A detailed method summarization is shown in Table 7. We highlight current limitations and suggest future research directions, providing valuable insights to advance PLLM development. Beyond taxonomy, we articulate a vision for memory-centric PLLMs, highlighting three desired capabilities—remember, adapt, and evolve—and analyze their inherent trade-offs among performance, privacy, and efficiency. We further examine domain-specific challenges and discuss future directions. Together, these contributions not only summarize the state of the art but also chart promising pathways for advancing PLLM research and practical deployment.
Appendix A Notation Summary
| Symbol | Description |
| • Spaces | |
| Space of all personalized data | |
| Latent abstraction space (preferences, profiles, tendencies) | |
| Personalization state space (style, policy, reasoning priors, etc.) | |
| Predefined human preference dimensions (e.g., “helpfulness”) | |
| Query space | |
| Output / response space | |
| General or world knowledge available to the LLM | |
| Set of users | |
| • Functions / Mappings | |
| Abstraction / summarization mapping | |
| Personalization encoder (possibly query-aware) | |
| Extraction operator (factual lookup) | |
| Abstraction operator (summary-based answering) | |
| Conditional distribution for generalization queries | |
| • Models | |
| Base large language model (non-personalized) | |
| Personalized model for user | |
| Personalization operator (e.g., prompting, adaptation, alignment) | |
| Tunable parameters (within LLM or added modules, e.g., PEFT/LoRA) | |
| • Operations / Sets | |
| Explicit factual tuples contained in (for Extraction) | |
| • Variables | |
| The -th user | |
| Personalized data of user | |
| Query set of user | |
| The -th query of user | |
| Predicted response for (model output) | |
| Desired personalized output for (ground truth) | |
| Training pair: context/content and reference response | |
| Reference response associated with in training data | |
| Abstracted profile of user | |
| Personalized state for user under query | |